### v2

# A dimension reduction approach to edge weight estimation for use in spatial models

Abstract
Models for areal data are traditionally defined using the neighborhood structure of the regions on which data are observed. The unweighted adjacency matrix of a graph is commonly used to characterize the relationships between locations, resulting in the implicit assumption that all pairs of neighboring regions interact similarly, an assumption which may not be true in practice. It has been shown that more complex spatial relationships between graph nodes may be represented when edge weights are allowed to vary. Christensen and Hoff, (2024) introduced a covariance model for data observed on graphs which is more flexible than traditional alternatives, parameterizing covariance as a function of an unknown edge weights matrix. One potential issue with their approach is that each edge weight is treated as a unique parameter, resulting in increasingly challenging parameter estimation as graph size increases. Within this article we propose a framework for estimating edge weight matrices, reducing their effective dimension via a basis function representation. By further leveraging fast matrix calculations for derivative expressions, we enhance the performance and flexibility of covariance models parameterized by such matrices, and demonstrate the utility of our method in a series of illustrations, simulations and data examples.
Keywords: Areal data, covariance, line graph, eigenvector basis, CAR model, spatial deformation
1 Introduction
Studies involving areal, or region indexed data are common in many settings including ecology (e.g. Hanks and Hooten,, 2013), economics (e.g. Arbia,, 2012), public health (e.g. Jin et al.,, 2005) and sociology (e.g. Garner and Raudenbush,, 1991). Models for spatially indexed data are typically defined such that observations close to one another in space exhibit higher correlation than observations which are far apart Stein et al., (2004). In contrast to spatial models for point-indexed data, which tend to characterize between-location dependence as a function of the geographic distances between observations, models for areal data generally define spatial dependence in terms of neighborhood or adjacency structure (Cressie,, 1993).
The spatial structure of areal data may be represented using a graph, with each spatial region as a graph node and with edges existing between each pair of nodes corresponding to physically adjacent regions. First introduced by Besag, (1974), the most commonly used spatial model for areal data is the conditional autoregressive (CAR) model. Given vertex set $V$ and edge set $E$, we define the graph $G=(V,E)$ with $p=|V|$ nodes. We observe the random vector y, where $y_{1},...,y_{p}$ are observations at each node of $G$. A CAR model for y defines a Markov random field where the conditional distribution of each element of y is determined by the values of y at adjacent nodes:
$y_{i}|\textbf{y}_{-i}\sim N\left(\sum_{j\sim i}\alpha w_{ij}y_{j},\sigma^{2}\right),$ (1)
where $j\sim i$ indicates that node $j$ is adjacent to node $i$, $w_{ij}$ characterizes the strength of the relationship between nodes $i$ and $j$, $\sigma^{2}>0$ is a variance parameter, and $\alpha$ controls the degree of spatial correlation. Equation 1 results in a marginal distribution for y of
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\textbf{I}_{p}-\alpha\textbf{W})^{-1}),$ (2)
where W is a $p\times p$ adjacency or weights matrix defined such that $w_{ij}>0$ if $j\sim i$ and $w_{ij}=0$ otherwise. In order to produce a valid covariance matrix, $\alpha$ must be bounded between the largest and smallest eigenvalue of W (Ver Hoef et al.,, 2018). An alternative form of the CAR model is
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\text{diag}(\textbf{W}\textbf{1}_{p})-\kappa\textbf{W})^{-1}),$ (3)
[…]
While many variants of the CAR model and other similarly defined spatial autoregressive models have been developed and utilized, Ver Hoef et al., (2018) note that in practice almost all such models have been implemented using the unweighted adjacency matrix of $G$, with $w_{ij}=1$ if $i\sim j$, and $w_{ij}=0$ otherwise. The CAR model is frequently used to characterize spatial random effects or an “error process” within a larger generalized linear model (Besag et al.,, 1991); in many such contexts, interest in the covariance structure of the model has been secondary to inference on the fixed effects governing the model’s mean structure. However, the lack of a more complex weighting scheme results in a pattern of spatial dependence where the partial correlations between all pairs of adjacent regions are near-identical regardless of location within the overall spatial domain. Such models may be thought of as discrete space equivalents to stationary and isotropic covariance models for point indexed data in continuous space, properties which are frequently inappropriate for real world data (Guttorp et al.,, 1994). Despite likely under-utilization in practice, more complex weighting schemes for CAR models have been shown to be useful, as in Wall, (2004) who examines the implied correlation structures in various models for areal data, and Hanks and Hooten, (2013) and Ejigu and Wencheko, (2020) who define edge weights using linear combinations of environmental covariates. Weighting schemes with even greater flexibility are also possible and they potentially enable more realistic interaction structures in the discretely observed spatial processes.
Christensen and Hoff, (2024) introduced a covariance model for areal data exhibiting greater flexibility than the traditional CAR model. We refer to this as the graph deformation (GDEF) model. Given the graph $G=(V,E)$, they define $\Sigma=\text{Cov}(\textbf{y})$ as follows:
$\Sigma_{ij}$ $=\sigma^{2}\rho_{\nu}(d_{ij}),$ (4)
$d_{ij}$ $=\sqrt{(\textbf{e}_{i}-\textbf{e}_{j})^{\top}\{\textbf{L}^{+}\}^{2}(\textbf{e}_{i}-\textbf{e}_{j})},$
L $=\text{diag}(\textbf{W}\textbf{1}_{p})-\textbf{W},$
$\sigma^{2}$ $>0,\;w_{ij}>0\text{ if }i\sim j,\text{ else }w_{ij}=0,\text{ and }w_{ij}=w_{ji}.$
Here $\sigma^{2}$ is a scale parameter and $\rho_{\nu}(\cdot)$ is the Matérn correlation function (Matérn,, 1960) with smoothness parameter $\nu$. The term $d_{ij}$ may be thought of as the latent distance between nodes $i$ and $j$ and is defined as a function of unknown edge weights matrix W. The matrix $\{\textbf{L}^{+}\}^{2}=\textbf{L}^{+}\textbf{L}^{+}$ is the square of the generalized inverse of the graph Laplacian L, and $\textbf{e}_{i}$, $i=1,\ldots,p$ are the standard basis vectors. The edge weights $\{w_{ij}\}_{i\sim j}$ are estimated from the observed data and may be interpreted as the strength of connectivity between adjacent nodes or regions. Related models have been suggested for resistance distances in physics and chemistry (Klein and Randić,, 1993).
The GDEF model described in Equation 4 results in a valid covariance matrix for any combination of positive edge weights and the matrix W is an identifiable parameter. While Christensen and Hoff, (2024) demonstrate the flexibility and interpretive value of the GDEF approach, they acknowledge that the model has limited scalability when applied to large graphs and using their Markov chain Monte Carlo (MCMC) algorithm for estimation.
A key question for both the implementation of the GDEF model and within the literature concerning more complex weighting schemes for spatial autoregressive models is how to define and estimate the edge weights matrix W. The individual estimation of each unique edge weight, as in Christensen and Hoff, (2024), proves to be computationally impractical for graphs with more than a few hundred nodes, while also introducing potential identifiability issues for models that have not been carefully specified. On the other hand, the covariance structures resulting from previously proposed edge weighting schemes, such as defining edge weights as linear combinations of covariates, or scaling edge weights to be proportional to distances between region centroids or the length of the border shared by two regions, may not be flexible enough to accurately characterize some spatial processes.
The principle contributions from our work within this article are as follows:
  • We introduce a dimension reduction approach for the specification and estimation of an unknown edge weights matrix that utilizes eigenvectors from the Laplacian matrix of a network’s line graph as a basis. This method allows for the specification of CAR and other autoregressive models that are considerably more flexible than existing variants, while significantly improving the practical usability of the GDEF framework.
  • By embedding the basis function representation of W within a covariance model rather than directly defining spatial random effects as linear combinations of basis functions as has traditionally been done, we leverage the computational advantages of such approaches while producing a class of models with fewer restrictions on the imposed correlation structure.
  • By incorporating our framework for edge weight estimation within the GDEF framework, we produce a nonstationary spatial covariance model that may be interpreted through the lens of the spatial deformation approach of Sampson and Guttorp, (1992) that may be fit with no replicate observations of the spatial process.
In Section 2 of the article we provide a brief overview of necessary background information related to spatial basis functions, graphs, and spectral theory. In Section 3 we present our approach to edge weight estimation and illustrate some of its properties when incorporated into existing methods. In Section 4 we illustrate and evaluate the performance of our approach through a series of simulation studies and a real data example. We conclude the article with a discussion of potential applications and extensions.
2 Background
Given a simple, undirected graph $G=(V,E)$, random variables $y_{i}$, $i\in V$, and defining $\Sigma_{ij}=\text{Cov}(y_{i},y_{j})$, we parameterize the covariance matrix $\Sigma$ as a function of W, a matrix of unknown positive edge weights. We define $\mathcal{W}_{G}$ to be the space of all edge weight matrices W that may be defined given $G$ such that $w_{ij}>0$ if an edge exists between nodes $i$ and $j$ and $w_{ij}=0$ otherwise. Note that this means that when we speak of estimating W, we are referring only to the estimation of the edge weights within a fixed network structure, rather than trying to determine the existence or nonexistence of edges between observed nodes. As the number of nodes in $G$ increases, so too does the size of $\mathcal{W}_{G}$. To address this challenge, we here restrict the parameter space to $\textbf{W}\in\mathcal{W}^{B}_{G}\subset\mathcal{W}_{G}$, where $\mathcal{W}^{B}_{G}$ contains all possible edge weights matrices that can be obtained from a basis function representation of the non-zero edge weights.
2.1 Basis functions and spatial models
[…]
$y(\textbf{s})=\textbf{x}^{\top}(\textbf{s})\beta+z(\textbf{s})+\epsilon(\textbf{s}),$ (5)
[…]
Let $\textbf{z}=(z(\textbf{s}_{1}),...,z(\textbf{s}_{p}))$ denote the vector of spatial random effects at some collection of $p$ locations. Instead of modeling $\textbf{z}\sim N_{p}(\textbf{0},\Sigma)$ it is common to define $\textbf{z}=\textbf{M}\eta$, where M is an $p\times k$ matrix of $k\ll p$ basis functions, and $\eta\sim N_{k}(\textbf{0},\Phi)$. Reduced-rank statistical models representing spatial random effects using basis functions are widespread. Examples of this include, but are not limited to the kernel convolution approach (Higdon,, 1998), predictive processes (Banerjee et al.,, 2008), spatial splines (Sangalli et al.,, 2013), the stochastic partial differential equation approach (Lindgren et al.,, 2011), and various eigenvector and principle component based methods (Higdon et al.,, 2008; Hughes and Haran,, 2013; Lee and Haran,, 2022). Note that the majority of these methods were developed for point indexed data, though the framework of Hughes and Haran, (2013) is designed specifically for areal data. Applying methods designed for one data type to another is possible, but requires assumptions that may not be appropriate in all settings (Gelfand et al.,, 2001).
Basis function methods tend to be very computationally efficient relative to full-rank Gaussian process models. Notably, some approaches to basis functions specification are also capable of addressing issues such as nonstationarity and anisotropy (Schmidt and Guttorp,, 2020). It has often been observed that methods based on low-rank representations of spatial random effects may result in oversmoothing as well as bias in the estimation of variance components (Banerjee et al.,, 2010; Stein,, 2014; Datta et al.,, 2016). Additionally, when pre-specifying the design matrix M using some set of basis functions, the covariance of u is restricted to matrices of the form $\textbf{M}\Phi\textbf{M}^{\top}$: in many instances, $\Phi$ is assumed to be a diagonal matrix or multiple of the identity matrix, resulting in strong and potentially inappropriate assumptions about model covariance structure. This may be especially true if analysis goals go beyond spatial prediction and smoothing and include inference regarding model covariance. Rather than modeling spatial random effects for areal data as a linear combination of basis functions, we propose using basis functions in a novel parameterization of edge weights matrix W which is then incorporated into a CAR or GDEF model, resulting in a flexible class of full-rank covariance matrices that are less computationally demanding than the implementation used by Christensen and Hoff, (2024).
2.2 Line graphs and an eigenvector basis
For a graph $G$, one may define its line graph $L(G)$ as a graph with $q$ nodes corresponding to each of the edges of $G$, and edges between each pair of nodes corresponding to coincident edges in the original graph $G$. (Edges are considered coincident if they share a common node). Figure 1 illustrates the relationship between a graph $G$ and its line graph $L(G)$ for a graph containing $p=5$ nodes and $q=8$ edges.
[…]
Figure 1: The graph $G$ has corresponding line graph $L(G)$. The five nodes are $v_{1},\ldots,v_{5}$. The eight edges are $e_{1},\ldots,e_{8}$.
Let $\textbf{A}^{L}$ be the $q\times q$ adjacency matrix of $L(G)$, where $a^{L}_{ij}=1$ if nodes $i$ and $j$ of $L(G)$ are adjacent and $a^{L}_{ij}=0$ otherwise. Let $\textbf{L}^{L}=\text{diag}(\textbf{A}^{L}\textbf{1}_{q})-\textbf{A}^{L}$ be the Laplacian matrix of $L(G)$. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}$ is the eigendecomposition of $\textbf{L}^{L}$, the columns vectors of V form an orthonormal basis for $\mathbb{R}^{q}$. The eigenvalues of Laplacian matrices are non-negative and generally ordered from smallest to largest, with smallest non-zero eigenvalues corresponding to the most “important” eigenvectors. Laplacian matrices for connected graphs are positive-semidefinite and have exactly one eigenvalue (denoted $\lambda_{1}$) equal to 0 with corresponding eigenvector $\textbf{v}_{1}\propto\textbf{1}$. Despite some referring to $\textbf{v}_{1}$ as the “trivial” eigenvector, it has interpretive value as an intercept column within an eigenvector basis. Considerable work has been done in the field of spectral graph theory investigating the properties of the eigenvalues and eigenvectors of Laplacian matrices (Spielman,, 2012).
Laplacian matrices have many applications in various types of data analysis. Within the field of signal processing for data on undirected graphs, the matrix $\textbf{V}^{-1}$ defines the graph Fourier transform, and is used to decompose signals in terms of the eigenvectors of the Laplacian (Sandryhaila and Moura,, 2013; Ortega et al.,, 2018). In the field of spatial statistics, the eigenvectors of graph Laplacians have been used as a basis to define a smoothly varying processes over networks, as in Upton and Carvalho, (2024) who model urban crime on a street network by estimating node-level regression coefficients in such a manner. One of the most novel contributions of our work in this article is the utilization of the eigenvectors of the Laplacian matrix of a network’s line graph to define a process over its edges rather than its nodes. Similar methods were applied recently in Josephs and Upton, (2025), who estimated the efficacy of different NBA lineups by utilizing a hypergraph representation of all players and player groupings, though the nature of the networks they analyzed and the random variables modeled differ considerably from those in focus here.
In order to understand how the eigendecomposition of the line graph Laplacian (LGL) provides a useful basis for modeling the edge weights within a network, we visualize the the original graph with edges weighted by the the corresponding elements of the LGL eigenvectors. Figure 2 depicts a selection of LGL eigenvectors for a $\times 30$ 30 lattice graph, while Figure 3 shows that of the 100 counties of North Carolina. Note the harmonic behavior of these eigenvectors, with later eigenvectors exhibiting higher frequency oscillations between positively and negatively valued regions of the graph. This makes selection of the first $k$ eigenvectors (corresponding to the $k$ smallest eigenvalues) a natural choice of basis, with the “most important” initial eigenvectors capturing larger-scale trends in edge weight behavior.
Figure 2: A selection of LGL eigenvectors for the $\times 30$ 30 lattice grid.
Figure 3: A selection of LGL eigenvectors for the graph of the 100 counties of North Carolina.
3 Method
Equations 3 and 4 describe two models for areal data, each parameterized by unknown edge weights matrix W. Let w be the $q$-length vector containing the unknown weights associated with the edges of graph $G$. The $p\times p$ edge weights matrix W may be constructed from w and vice versa. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}=\textbf{L}^{L}$ is the eigendecomposition of the LGL, let $\textbf{V}^{k}$ be the $q\times k$ matrix containing the $k$ eigenvectors of $\textbf{L}^{L}$ associated with its $k$ smallest eigenvalues. We can now represent w as follows:
w $=\text{exp}(\textbf{V}^{k}\eta),\;\eta\in\mathbb{R}^{k},$ (6)
where $\eta$ is a $k$-length vector of basis coefficients that fully characterizes the $p\times p$ edge weights matrix $\textbf{W}(\eta)$. For all $\eta\in\mathbb{R}^{k}$, (and $\sigma^{2}>0$ and $|\kappa|<1$) the CAR covariance matrix $\Sigma=\sigma^{2}(\text{diag}(\textbf{W}(\eta)\textbf{1}_{p})-\kappa\textbf{W}(\eta))^{-1}$ and the GDEF covariance matrix defined in Equation 4 are positive definite matrices.
A discussion and comparison of the properties of the CAR and GDEF modeling approaches is provided in the supplemental material. In general we find that CAR model is most naturally suited for applications in which the primary focus is the estimation of a dataset’s mean structure; in such settings the CAR framework provides an efficient way to account for spatial random effects that may otherwise obstruct inference on the fixed effects. When inference regarding a dataset’s covariance structure is an important modeling consideration, many authors have noted that the CAR covariance (even when utilizing weighting schemes) has properties that are undesirable or restrictive in many scenarios. See for example Wall, (2004); Besag and Mondal, (2005); Paciorek, (2013). While we find that incorporating the edge weighting scheme of Equation 6 into CAR models improves their flexibility, the primary focus within the remainder of this article will be on the use of the LGL eigenvector basis within the GDEF model.
3.1 Visualizations of method and interpretation of basis coefficients
As discussed in Christensen and Hoff, (2024) regarding the GDEF model, each possible edge weights matrix W corresponds to an embedding of the graph in high-dimensional Euclidean space which is unique up to isometry. The distances between nodes in this embedding may be thought of as the “intrinsic” distances between regions. The GDEF model may be viewed as the graph equivalent of the spatial deformation approach for modeling nonstationary data defined by Sampson and Guttorp, (1992), where the deformation function is obtained by estimating the edge weights. This interpretation leads us to say that nodes in regions containing high valued edges are “closer” to each other than regions with low valued edges. While the CAR model does not directly admit a distance based interpretation of W, it can be said of both the CAR and GDEF approaches that edges with large weights correspond to greater connectivity between incident nodes.
Let $\textbf{W}(\eta)$ be the $p\times p$ matrix of edge weights parameterized by $\eta\in\mathbb{R}^{k}$. Equation 6 states that the log of the $q$-length vector containing the unique edge weights in $\textbf{W}(\eta)$ is a linear combination of a collection of $k$ basis functions of length $q$, with $\textbf{V}^{k}$ defined as the $k$ eigenvectors of the LGL corresponding to its $k$ smallest eigenvalues. We wish to provide some insight into the interpretation of the coefficient vector $\eta=(\eta_{1},...,\eta_{k})^{\prime}$ and illustrate some of the covariance patterns this construction can produce.
Recall that $\eta_{1}\textbf{v}^{k}_{1}$ is constant and may be viewed as the intercept for the log edge weights. Because $\text{exp}(\textbf{V}^{k}\eta)=\text{exp}(\eta_{1}\textbf{v}^{k}_{1})\odot\text{exp}(\textbf{V}^{k}_{-1}\eta_{-1})$, the vector $\text{exp}(\eta_{1}\textbf{v}^{k}_{1})$ may be viewed as a scaling factor for the edge weights matrix $\textbf{W}(\eta_{-1})$ which is defined using the other $k-1$ basis functions. In the GDEF model, this scaling factor is directly related to the range of spatial correlation, while in the CAR model, the scaling factor is redundant when $\sigma^{2}$ is estimated. As such, we recommend fixing $\eta_{1}=0$ when using the basis function approach to characterizing the edge weights matrix for CAR models. Subsequent coefficients in $\eta$ may be interpreted in context of the eigenvectors with which they are associated. As depicted in Figure 2, the second LGL eigenvector of a $\times 30$ 30 grid characterizes a contrast between the top and bottom of the spatial domain. Figure 4 depicts samples from a GDEF covariance model defined using only the first two LGL eigenvectors and smoothness parameter $\nu=3/2$. Across samples it can be seen how observations in the top of the spatial domain are highly correlated, while the effective range of spatial correlation is much smaller on the opposite side. Spatial covariance patterns such as this do appear in real world settings; for example Paciorek and Schervish, (2006) present a data set for precipitation in Colorado in which observations east of the Rocky mountains are highly correlated over long distances, whereas spatial correlations between observations in the mountainous western half of the state decay rapidly over short distances.
Figure 4: Samples from $N(\textbf{0},\Sigma)$ over a $\times 30$ 30 grid using two basis functions with $\eta_{1}=\eta_{2}=50$. Note how spatial correlation decays at different rates across the spatial domain.
Different linear combinations of basis functions can produce a wide array of covariance structures. Figure 5 depicts a linear combination of four eigenvectors that results in covariance where observations at the borders of the spatial domain are highly correlated, while observations in the interior are not. In effect locations at opposite corners of the grid are “closer” together than locations that are both near the center. The more basis functions are included, the more flexible the covariance, though there are generally diminishing returns as the number of eigenvectors used increases due to the higher spatial frequencies exhibited by later eigenvectors. Omitting these high frequency eigenvectors essentially assumes that the edge weights themselves exhibit some degree of spatial smoothness. The question of how to select the number of basis functions is explored later in Section 4, and a discussion regarding the potential inclusion of environmental covariates within our model for edge weights is provided in the supplemental material.
Figure 5: Illustration of how limited number of basis functions may combine to produce complex covariance structures. $\Sigma$ defined using GDEF covariance with $\nu=3/2$.
3.2 Parameter estimation and model fitting
Much of the spatial deformation literature assumes settings in which there are replicate observations of the spatial process whose covariance is being estimated. Commentary on this assumption is provided in the supplemental material. Should we have replicate observations, we record them as $\textbf{y}_{t}\in\mathbb{R}^{p}$ for $t\in 1,...,n$. These observations may be stored in the $n\times p$ data matrix Y, where $y_{ti}$ is the $t$th observation at the $i$th node of graph $G$. For now, we assume that $\textbf{y}_{t}\overset{iid}{\sim}N_{p}(\textbf{0},\Sigma)$, with $\Sigma$ parameterized by $\theta=(\sigma^{2},\eta)$ for the GDEF model and $\theta=(\sigma^{2},\eta,\kappa)$ for the CAR model. Christensen and Hoff, (2024) only considered settings with multiple observations of the spatial process. This was necessary due to each edge weight being a unique parameter and the fact $q>p$ in almost all settings. Importantly, the dimension reduction of the basis function approach to estimation of W allows us to model the covariance of spatial processes with only a single realization ($n=1$).
The maximum likelihood estimator (MLE) for $\theta$ may be obtained using Fisher scoring which requires computation of the first and second derivatives of the log likelihood $\ell(\theta;\textbf{Y})=-\frac{n}{2}\log|\mbox{\boldmath$\Sigma$}(\mbox{\boldmath$\theta$})|-\frac{1}{2}\sum_{t=1}^{n}\textbf{y}^{T}_{t}\mbox{\boldmath$\Sigma$}^{-1}(\mbox{\boldmath$\theta$})\textbf{y}_{t}$ with respect to $\theta$. Given data Y and an initial set of parameter values $\theta^{0}$ we can maximize the likelihood by iteratively calculating
$\theta^{l}=\theta^{l-1}-\gamma\;\mathbb{E}\left[\textbf{H}_{\ell}(\theta^{l-1})\right]^{-1}\left(\nabla_{\ell}(\theta^{l-1})\right),\hskip 14.22636ptl=1,2,\ldots,$ (7)
until $\theta^{l}$ converges to the MLE. Here $\gamma$ is a scaling parameter between 0 and 111. Smaller $\gamma$ provides more robust optimization, while $\gamma=1$ will generally be faster. Identifying the impact of individual basis function coefficients on the likelihood is fairly challenging, but details on the efficient matrix forms of $\nabla_{\ell}(\theta)=\frac{\partial\ell(\theta;\textbf{Y})}{\partial\theta}$ and $\textbf{H}_{\ell}(\theta)=\frac{\partial^{2}\ell(\theta;\textbf{Y})}{\partial\theta^{2}}$ are provided in the Supplemental material.
Uncertainty quantification may be conducted using the approximation
$\hat{\theta}\sim N\left(\theta,-\left[\textbf{H}_{\ell}(\hat{\theta})\right]^{-1}\right).$ (8)
Note when employing the algorithm in Equation 7 that we use the expectation of the Hessian matrix $\textbf{H}_{\ell}(\theta)$, which is less computationally intensive to obtain than the Hessian matrix itself. We only need to compute $\textbf{H}_{\ell}(\theta)$ once after obtaining the MLE in order to use the approximation in Equation 8.
3.3 Prediction
The correlation structures of most models for areal data, including the GDEF and especially the CAR models, are sensitive to the structure of the graph on which they are defined (Wall,, 2004; Christensen and Hoff,, 2024). As such, prediction for new or unobserved locations has to be approached more carefully than when using models for point-indexed data to ensure validity.
The log likelihood of $\textbf{y}_{t}$ and its derivatives with respect to $\theta$ are still defined when $\textbf{y}_{t}$ is not fully observed at all nodes within a graph. If one wishes to make predictions for unobserved nodes or new locations, one simply needs to ensure that the graph $G$ used to fit the model is specified such that its nodes and edges are defined for all relevant locations, both where data are observed and where one wishes to predict. Because the LGL basis functions used to characterize the edge weights matrix W are defined over the full network rather than locally, edge structure in unobserved regions of the graph is still estimable. After obtaining estimates for all parameters needed to define $\Sigma(\theta)$, it is straightforward to use the conditional distribution of $\textbf{y}_{\text{predicted}}|\textbf{y}_{\text{observed}}$ to obtain predictions and prediction intervals.
4 Illustrations
In this section we evaluate the performance of the GDEF model using the LGL eigenvector basis in both simulated and real-world data settings. We begin with two simulation studies. In the first, we consider how to select the number of basis functions when implementing our method. The second simulation assesses model performance under misspecification, evaluating the Kullback-Leibler divergence between the true data generating distribution and the estimated distribution. The Supplemental material contains an additional simulation study demonstrating coverage rates of the approximate confidence intervals obtained using Equation 8, as well as additional details on the two simulations presented here.
4.1 Simulation 1: Number of basis functions
We defined $\times 10$ 10 and $\times 20$ 20 lattices obtaining LGL eigenvectors for each. Instead of simulating edge weights directly from Equation 6, we generate an edge weights matrix W from an ICAR model using the line graph adjacency matrix to define the associations between edges. Figure 3 of the supplemental material for this article depicts one such sampled set of edge weights over a $\times 20$ 20 lattice. The spatial dependence in the simulated edge weights themselves means that W can be well approximated using our method. Larger numbers of basis functions will improve this approximation, but at the cost of model complexity and potential computational instability. The simulated W was used to define a covariance matrix according to Equation 4 with $\nu=3/2$ and $\sigma^{2}=1$. Independent samples were generated from $N_{p}(\textbf{0},\Sigma)$ and stored in $n\times p$ matrix Y for $n\in\{1,10,50\}$. For each simulated Y we fit a GDEF model using Matérn covariance with $\nu=3/2$ using $k\in\{10,20,...,100\}$ basis functions and evaluated AIC and BIC for each of the 10 model fits. The simulation was repeated 10 times. Table 1 contains information regarding the AIC- and BIC-preferred choices of $k$.
Mode $k$ (AIC) Mean $k$ (AIC) Mode $k$ (BIC) Mean $k$ (BIC)
$p=100,\;n=1$ 10 12 10 10
$p=100,\;n=10$ 40 58 10 11
$p=100,\;n=50$ 100 100 40 42
$p=400,\;n=1$ 20 37 10 10
$p=400,\;n=10$ 100 99 30 39
$p=400,\;n=50$ 100 100 100 99
Table 1: Mode $k$ indicates the most commonly selected number of basis functions, while Mean $k$ indicates the average number of selected basis functions. Larger grids and larger sample sizes result in preference for higher $k$.
Across simulation settings, we see that models with higher $k$ perform better on the larger grid and with bigger sample sizes. Using the information in Table 1, a potential starting point is to select $k=\sqrt{np}$ when the model reasonably approximates the true data generating process, as was the case within this simulation. Optimal choice of $k$ will also be contingent on the nature of the spatial process itself, as evidence of more complicated patterns of spatial dependence may indicate that larger numbers of basis functions should be used.
4.2 Simulation 2: Model performance under misspecification
We compare the performance of the GDEF model with others in a setting where the GDEF model is misspecified. Data is generated by the nonstationary spatial process obtained by the spatial deformation approach of Sampson and Guttorp, (1992). This defines the process covariance
$\text{Cov}(z(\textbf{s}),z(\textbf{s}^{\prime}))=\sigma^{2}\rho(||g(\textbf{s})-g(\textbf{s}^{\prime})||),$ (9)
where $g(\cdot)$ is a transformation from the geographic location s to its location $g(\textbf{s})$ in a “deformed” space wherein covariance is stationary and isotropic. Data are simulated by first defining a $\times 15$ 15 lattice. We then generate random deformations of the lattice by contracting or expanding the grid towards or away from randomly sampled locations. Let $\textbf{D}^{D}$ be the $\times 225$ 225 matrix of Euclidean distances between the nodes of the deformed lattice. Applying a Matérn correlation function to $\textbf{D}^{D}$ results in a covariance matrix that is nonstationary with respect to D, the matrix of Euclidean distances between the original grid point locations. We generate $n$ observations from the following model
$\textbf{y}_{t}$ $=\beta_{0}\textbf{1}+\textbf{z}_{t}+\epsilon_{t},$ (10)
$\textbf{z}_{t}$ $\sim N_{p}(\textbf{0},\sigma^{2}\rho_{5/2}),$
$\epsilon_{t}$ $\sim N_{p}(\textbf{0},\tau^{2}\textbf{I}_{p}),$
where $\textbf{z}_{t}$ is a vector of spatial random effects with correlation matrix $\rho_{5/2}$ defined using a Matérn correlation function with smoothness $\nu=5/2$. We set $\beta_{0}=0$, $\sigma^{2}=0.9$, $\tau^{2}=0.1$ and a range parameter of 333. We fit the GDEF and other models to the simulated data Y as illustrated by Figure 6, which depicts a deformed lattice, samples of $\textbf{y}_{t}$ produced by that deformation, and an estimated edge weights matrix.
Figure 6: Distances between deformed grid locations define a nonstationary covariance matrix $\Sigma$. Data Y generated from $N_{p}(\textbf{0},\Sigma)$. GDEF model is fit to Y, producing the depicted $\hat{\textbf{W}}$ and corresponding $\hat{\Sigma}^{GDEF}$.
We randomly generated 125 unique grid deformations. For each deformation we sampled Y as described above for $n\in\{1,2,3,5,10,25,50\}$ and fit Equation 10 while defining the covariance of the random effects $\textbf{z}_{t}$ using the following five submodels: a GDEF model with $\nu=3/2$ and $k=15$ basis functions, the spatial deformation approach of Sampson and Guttorp, (1992) (S&G) as implemented by the deform R package (Youngman,, 2023), a simple Matérn covariance model with $\nu=5/2$ and using the Euclidean distances between non-deformed lattice nodes, the method presented in Hughes and Haran, (2013) (H&H) for modeling random effects in areal data using 50 basis functions, and a standard CAR model using an unweighted adjacency matrix. The spatial deformation approach and the stationary Matérn covariance model treat the data as though it were point-indexed, while all other approaches utilize only the lattice’s graphical structure; note however, that the true data generating process exists in continuous space.
Let $\Sigma=\sigma^{2}\Phi+\tau^{2}\textbf{I}_{p}$ be the marginal covariance of $\textbf{y}_{i}$ integrated over the random effects $\textbf{z}_{i}$. For each Y and fitted model, we obtain the MLE $\hat{\Sigma}$ and compute $\text{KL}(N_{p}(\textbf{0},\Sigma),N_{p}(\textbf{0},\hat{\Sigma}))$, the Kullback-Leibler (KL) divergence between the true data generating distribution and the distribution characterized by the estimated covariance. As both distributions are mean-zero and Gaussian, KL divergence may be computed using only the covariance matrices $\Sigma$ and $\hat{\Sigma}$:
$\text{KL}(\Sigma,\hat{\Sigma})=\frac{1}{2}\left(\text{log}\frac{|\hat{\Sigma}|}{|\Sigma|}+\text{tr}\left(\hat{\Sigma}^{-1}\Sigma)\right)-p\right).$ (11)
Table 2 depicts the average KL divergence between the true distributions and the distributions estimated by the five models. Note that the true distribution is different for each of the 125 replicates which used a unique deformation. We also consider the frequency with which each model had the smallest KL divergence under the same settings. As can be seen, the GDEF model performs well across sample sizes, though the simple Matérn and CAR models frequently produce competitive estimates of $\Sigma$ when $n=1$. It is intuitive that simpler models would perform better in contexts with limited data available, and visual inspection of simulation output suggests that the simple Matérn and CAR models perform best in comparison to the GDEF approach when the randomly generated deformation exhibits less dramatic patterns of warping. The GDEF model is the best performing in settings with 10 or fewer replicates, though in settings with $n=25$ and $n=50$ it is surpassed on average by the S&G approach, which under the settings of this simulation study is closest to the true data generating model. The fact that the GDEF model consistently outperforms S&G for $n\leq 10$ suggests that its internal structure provides some degree of stability and regularization when estimating complex covariance structures using limited numbers of replicates.
GDEF S&G Matérn 5/2 H&H Std. CAR
$n=1$ 42.54 (0.552) - 45.22 (0.24) 70.85 43.4 (0.208)
$n=2$ 28.2 (0.928) 73.6 42.35 (0.04) 69.35 42.28 (0.032)
$n=3$ 24.24 (0.984) 40.29 40.87 (0.016) 75.78 41.88
$n=5$ 21.93 (0.992) 32.33 (0.008) 39.75 58.58 41.67
$n=10$ 20.6 (0.896) 26.54 (0.104) 39.13 51.75 41.56
$n=25$ 19.83 (0.376) 19.09 (0.624) 38.65 50.10 41.44
$n=50$ 19.55 (0.192) 17.02 (0.808) 38.36 49.7 41.39
Table 2: Average $\text{KL}(\Sigma,\hat{\Sigma})$ for each model and number of samples. In parentheses is the frequency across simulations that a given model produced the smallest KL divergence for a sample of size $n$. The GDEF model performs well across settings, and especially with limited numbers of replicates.
Note that the spatial deformation approach (S&G) cannot be fit for $n=1$. This highlights one of the key advantages of our approach. Much of the literature regarding the spatial deformation framework requires that the spatial process be observed multiple times, and in some implementations requires that $n>p$. However, it is extremely common for spatial analyses to be conducted on datasets with little or no replication. Whether the deformation function is estimated using thin plate splines as in Sampson and Guttorp, (1992) or a Gaussian process prior as in Schmidt and O’Hagan, (2003), the inherent flexibility of the framework requires considerable amounts of data (or strong prior assumptions) to avoid over-fitting. As seen here the GDEF model using the basis function representation of W as presented in this paper can effectively estimate complicated patterns of spatial dependence for areal data and performs well in a setting where the true spatial process exists in continuous rather than discrete space.
4.3 Example: Mercer and Hall wheat yield data
We conclude this section with an analysis of the classic wheat yield dataset from Mercer and Hall, (1911) and which is available in the spData R package. The version of the data used in the package was taken from Cressie, (1993). Mercer and Hall, (1911) considered a one acre parcel of farmland that was divided into 500 rectangular plots and planted with wheat. Each plot received the same treatment and soil quality was perceived to be essentially uniform at the time of planting. At harvest, the yield of each plot in pounds of grain was recorded. Figure 7 depicts the yields for each plot, which were arranged in 20 rows each containing 25 plots. Mercer and Hall discuss the variability present in the data and explore the possibility of location-based effects on yield. Noting evidence of spatial correlation, the wheat yield data was analyzed in Whittle, (1954), which first introduced the simultaneous autoregressive (SAR) model, and in Besag, (1974), which first introduced the CAR model. The authors of both papers found the fit of their models to the wheat yield data unsatisfactory. Columns (4,7 and 10) in Figure 7 have larger average yields than others and the within-column correlation appears to be stronger than within-row correlation. By modeling the data using fixed row and column effects (Cressie,, 1993) found that a stationary and isotropic kriging model with plot locations given by centroids resulted in satisfactory fit, though it is worth considering whether the inclusion of row and column effects in a model for this data provide interpretive value or merely serve to improve the fit of an otherwise poorly specified model.
Figure 7: The Mercer Hall wheat yield data. Each cell corresponds to an agricultural plot; response variable is pounds of grain at harvest.
[…]
Figure 8: Estimated log edge weights (left) for the Mercer-Hall data which in turn define the covariance structure of z, the “de-noised” version of the centered data (right).
Est. 95% CI
$\eta_{1}^{row}$2.195 (-3.551, -0.838)
$\eta_{1}^{col}$ 1.308 (0.286, 2.329)
$\sigma^{2}$ 0.146 (0.101, 0.191)
$\tau^{2}$ 0.073 (0.027, 0.118)
$\beta_{0}$ 3.949 (3.871, 4.026)
Table 3: MLEs and confidence intervals for model parameters.
We can also obtain an estimate of z, which may be thought of as the smoothed or “de-noised” version of the centered data. Because y and z are each normally distributed, the conditional distribution of z is straightforward to obtain:
$\textbf{z}|-\sim N_{p}\left(\left(\Phi^{-1}/\sigma^{2}+\textbf{I}_{p}/\tau^{2}\right)^{-1}(\textbf{y}-\beta_{0}\textbf{1})/\tau^{2},\left(\Phi^{-1}/\sigma^{2}+\textbf{I}_{p}/\tau^{2}\right)^{-1}\right).$ (12)
We define $\hat{\textbf{z}}=\left(\hat{\Phi}^{-1}/\hat{\sigma^{2}}+\textbf{I}_{p}/\hat{\tau^{2}}\right)^{-1}(\textbf{y}-\hat{\beta_{0}}\textbf{1})/\hat{\tau^{2}}$ and provide a plot of $\hat{\textbf{z}}$ in the right half of Figure 8. Note that regions where the estimate of z are smoother correspond to larger estimated edge weights. Moran’s I (Moran,, 1950) is a test statistic indicating the presence or absence of spatial autocorrelation. Moran’s I for the estimated residuals of our model $\hat{\epsilon}=\textbf{y}-\hat{\beta_{0}}\textbf{1}-\hat{\textbf{z}}$ is 0.022, with a p-value of 0.460, indicating a lack of spatial correlation in $\hat{\epsilon}$. The Shapiro-Wilk normality test (Shapiro and Wilk,, 1965) on the estimated residuals results in a p-value of 0.262 indicating that the normality assumption for $\epsilon$ is not violated.
To assess the out-of-sample predictive performance of the GDEF model using the LGL eigenvector basis on this data set in comparison to existing methods, we perform ten-fold spatial cross validation, partitioning the 500 plot locations into 20 five-by-five blocks of plots, then splitting the 20 blocks into ten sets of two, resulting in a total of 50 plots per split. In addition to the GDEF model described above, we fit each of the following models to the test-training splits: a Matérn covariance model with $\nu=5/2$ and distances computed using plot centroids, the spatial random effects model of Hughes and Haran, (2013) using 50 basis functions, a standard CAR model, and a CAR model with two correlation parameters, one for each of the within-row and within-column dependencies, a modeling consideration similar in nature to the split we made of the intercept column in the basis representation of the GDEF edge weights. For each model and test-training split, we compute the root mean square error (RMSE) and mean absolute error (MAE), and produce 95% prediction intervals for all elements of the test set. In addition to our comparison of predictive performance, we wish to evaluate the appropriateness of each model for the wheat yield data using residual diagnostics. Each model was fit using the full data set and estimated residuals were obtained in a manner similar to the process described in the previous paragraph for the GDEF model. Moran’s I is computed for each model to assess whether the estimated residuals had any spatial dependence and the Shapiro-Wilk test is applied to assess the normality of the residuals. The left side of table 4 contains the cross-validated average RMSEs, MAEs, coverage rates, and prediction interval widths for each model, while the right side contains the p-values associated with the Moran’s and Shapiro-Wilk tests.
RMSE MAE Coverage PI Width Moran’s I Shapiro-Wilk
GDEF 0.410 0.331 0.904 1.450 $p=0.460$ $p=0.262$
Matérn 0.436 0.355 0.936 1.689 $p=0.000$ $p=0.028$
H&H 0.684 0.552 0.898 2.209 $p=0.000$ $p=0.934$
Std. CAR 0.421 0.340 0.940 1.688 $p=0.000$ $p=0.002$
CAR 2-param. 0.422 0.342 0.936 1.647 $p=0.000$ $p=0.008$
Table 4: Comparison of predictive metrics and results from assumption tests across model. In addition to having the smallest out-of-sample prediction error, the GDEF model is the only one that produces estimated residuals that exhibit no spatial correlation.
The GDEF model has the lowest prediction error among models considered. In addition to this, it is the only model for which the Moran’s I statistic results in a failure to reject the null hypothesis of no spatial dependence in the model’s residuals. In other words, the GDEF model is the only model among these options with a covariance structure that is capable of fully describing the spatial dependence pattern present in the wheat yield data. Among models compared, the GDEF model produced the narrowest prediction intervals, though the coverage rate of these intervals appears to be somewhat under target, suggesting that the estimated marginal variance of predictions from the GDEF model may be too low; further investigation of this could be warranted. Despite considerably greater model complexity than the alternatives, the GDEF model’s out-of-sample predictive performance suggests that it is not overfit to these data and represents an improvement over simpler, more standard methodologies. While it could be argued that the differences in predictive performance between the GDEF and other models is not dramatic, it seems clear that the underlying covariance structure of the wheat yield data is most fully captured by our proposed methodology. We find this to be notable given the dataset’s history and reputation as a challenging spatial modeling problem.
5 Discussion
Within this article we introduced a framework for estimating edge weight matrices using the eigenvectors of the line graph’s Laplacian matrix. Utilizing matrix calculations for the derivatives of the log likelihood with respect to basis function coefficients, we show that this method may be fit efficiently and how it can be used to enhance existing covariance models for areal data. We also evaluated properties of our approach in a series of simulations and examples that demonstrated its capacity to represent complex covariance structures while significantly improving the computational efficiency of the GDEF model introduced by Christensen and Hoff, (2024). We conclude this article with a brief discussion of potential avenues for future work and model extensions.
Part of our preference for using the GDEF model in conjunction with the LGL basis rather than the CAR model stems from findings that the latter characterizes fundamentally non-smooth spatial processes. Rue and Held, (2005) and Lindgren et al., (2011) demonstrate that it is possible to define Markov random fields (MRFs) that approximate smooth spatial processes by allowing some edge weights to be negative. The space of non-CAR MRFs is worth exploring due to their computational advantages. It may be possible to adapt our method for edge weight estimation to allow for the inclusion of negative weights in sparse precision matrix formulations, though additional work would be necessary to ensure that a valid covariance matrix is always produced.
While we have discussed estimation of W within the context of explicitly defining a covariance matrix, there are other settings in which a coherent framework for flexibly defining an edge weights matrix may be useful. For example, resistance distance is a graph metric constructed using a matrix of non-negative edge weights that has seen widespread use in many statistical and ecological applications (Klein and Randić,, 1993; Hanks and Hooten,, 2013; Dickson et al.,, 2019). Outside the context of distances and covariance, an edge weights matrix can be used to characterize the connectivity of different entities within any network.
An additional topic we would like to explore is how our method could be used to characterize the covariance of spatial processes on manifolds or restricted spatial domains. While most spatial models have been developed assuming that data are observed on some well-defined subset of $\mathbb{R}^{2}$, there is increased interest in the development of methods that can be applied to more complex domains in which Euclidean distances between observations do not accurately represent proximity with respect to the domain’s underlying geometry (Yang and Dunson,, 2016). Mathematical distinctions exist between settings in which data are observed in Euclidean space but concentrated near a lower-dimensional manifold such as the surface of a sphere and settings in which an $n$-dimensional spatial process exists only on an irregular subset of $\mathbb{R}^{n}$ due to underlying physical constraints (Dunson et al.,, 2022). Whether data is observed on a true manifold or a restricted domain, we could define a graph based on the set of observed locations, defining edges between locations based on neighbors within a small radius or through use of the Delauney triangulation (Hjelle and Dæhlen,, 2006). The LGL eigenvectors of this graph could then be utilized as in this article to define a valid covariance model over the complex spatial domain; furthermore, the flexibility of our edge weight estimation approach would ultimately allow covariance to be nonstationary with respect to the domain’s underlying geometry.
Supplemental Materials
File containing additional simulation results, details on computation, and further discussion of relevant topics. Supplemental materials available online.